A body starting from rest and moving with constant acceleration covers 20m in the first second. The distance travelled by it in 2ND second is :
1
Expert's answer
2018-12-12T09:58:08-0500

Answer on Question #83750, Mechanics/Relativity

A body starting from rest and moving with constant acceleration covers 20m in the first second. The distance travelled by it in 2ND second is :

Solution

x1 = v0 t + 1/2 (at2)

where x1 = 20m, v = 0, t= 1 s

Find a:

20 = 0*1 + 1/2 *(a*12)

a=40 m/s2

Find how far will the body move after time t= 2s

x2 = v0 t + 1/2 (at2)

where v = 0, t= 2 s, a = 40 m/s2

x2= 0*2 + 1/2 * (40*22)

x2= 80 m

Find the distance trevelled by a body in the 2ND second:

x= x2 - x1 = 80m -20 m = 60 m

Answer: 60 m